OpenStudy (anonymous):

You roll a pair of dice five times. Find the probability of each of the following outcomes. a) You get doubles exactly once b) You get exactly three sums of 7 c) You get at least one sum of 7 d) You get at most one sum of 7 I've been using the formula P(x=k)=(n / k)P^k (1-p)^n-k For a) I got 0.4019, not sure if that's correct I'm having trouble with the rest

OpenStudy (anonymous):

What distribution is that?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Ohhhh binomial distribution

OpenStudy (anonymous):

Yeah sorry it's the only one I've learned

OpenStudy (anonymous):

For the first one I have it set up like this n=5 success= doubles p=1/6 x= number of times getting doubles Then P(x=1)= (5 / 1 )(1/5)^1(1-1/6)^5-1 Where the (5 / 1) isn't supposed to have the / symbol but I've no idea how to show it's five over one without it.

OpenStudy (anonymous):

that's little p I'm not really sure how I'm supposed to do this :/

OpenStudy (anonymous):

wait lemme reread my notes here. I may be getting carried away

OpenStudy (anonymous):

X= number of times the success should happen. And since you want doubles only once for the first one, it's one. For b) it should be 3.

OpenStudy (anonymous):

That's what I think it means at least.

OpenStudy (anonymous):

See here is my issue. What is the probability of getting a double which in this case is our p

OpenStudy (anonymous):

How did you figure out what your p is

OpenStudy (anonymous):

1-1/6

OpenStudy (anonymous):

little p is the amount of times out of two six sided dice doubles could happen, which would be 1/6. Or 6/36

OpenStudy (anonymous):

This is how I'm reading my notes, not sure what you're thinking or reading

OpenStudy (anonymous):

Ahhhhhh now i seee itttttt

OpenStudy (anonymous):

Cool so I'm not wrong for once :p My problem is the next three questions.

OpenStudy (anonymous):

Ya but You are wrong in this case cuz p=1/6 so q=5/6

OpenStudy (anonymous):

q?

OpenStudy (anonymous):

I don't have a q in my formula

OpenStudy (anonymous):

P(x=1)= (5 / 1 )(1/5)^1(1-1/6)^5-1

OpenStudy (anonymous):

oh gotcha

OpenStudy (anonymous):

The real equation should be P(x=1)= (5 / 1 )(1/6)^1(1-1/6)^5-1

OpenStudy (anonymous):

I think I mistyped the equation in the first part, I have it written down correctly on my notes.

OpenStudy (anonymous):

which equals .4019 like you stated above

OpenStudy (anonymous):

Okk lets go onto the next one then

OpenStudy (anonymous):

Okay, so for this one I have P(> or = 3) So P(X=3)+P(X=4)+P(X=5)

OpenStudy (anonymous):

4+3=7 3+4=7 2+5=7 5+2=7 1+6=7 6+1=7 So there 6 ways we can get a sum of 7 when rolling to dies so p=1/6

OpenStudy (anonymous):

Yes, so for P(X=3) I have (5 / 3)(1/6)^3(1-1/6)^5-3

OpenStudy (anonymous):

Yup that looks correct

OpenStudy (anonymous):

But for (5 / 3) Is that equal to 1 also because it's in (n / n) format?

OpenStudy (anonymous):

(5*4*3*2*1)/((3*2*1)(2*1) 20/2=10

OpenStudy (anonymous):

I've never been shown that way. Why do you skip 4*3*2*1?

OpenStudy (anonymous):

Ok sorry just jumped a couple of steps cuz i thought u followed

OpenStudy (anonymous):

I'm sorry it's my first statistics class

OpenStudy (anonymous):

\[\frac{5!}{3!(5-3)!}\] \[\frac{5!}{3!*2!}\] \[\frac{5*4*3*2*1}{(3*2*1)(2*1)}\] \[\frac{5*4}{2*1}=\frac{20}{2}=10\]

OpenStudy (anonymous):

So that's what you do every time that you have a (5 / 3) or the likes? Not if there is a one involved?

OpenStudy (anonymous):

Yup here is the general formula